OpenStudy (anonymous):

Write the point-slope form of the equation passing through (5, -1) with a slope of 6.

OpenStudy (ranga):

Point slope form: y - y1 = m(x - x1) m = 6, x1 = 5, y1 = -1 y - (-1) = 6(x - 5) y + 1 = 6(x - 5) is the point slope form of the equation.

OpenStudy (anonymous):

Whoops! Copied that as a 2 instead of a 5, my bad

OpenStudy (anonymous):

(x0,y0)→(2,−1) Slope=m=6 Point Slope form: y−y0=m(x−x0) So: y+1=6(x−2) y+1=6x−12 And finally: y=6x−29

OpenStudy (ranga):

I think they want the answer in point-slope form and not slope-intercept form.

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &(5\quad ,&-1) \end{array} \\\quad \\ slope = m= 6\\ \quad \\ y-y_1=m(x-x_1)\) plug and chug =)